Please help if a=2, b=3, c=-4, d=-1 and e=0
(a/b)^2 - (c/d)^3
As I have no idea how to use fractions and exponents.
Thank you so very much

Answer by checkley71(8403)

(A/B)^2=(2/3)^2=(2^2)/(3^2)=4/9
(C/D)^3=(-4/-1)^3=(-4^3)/(-1)^3-64/-1=64
4/9-(64)=4/9-64=(4-64*9)/9=(4-576)/9=572/9=63 5/9 ANSWER.